OpenStudy (wesdg1978):

g(x)=1/(4th root(x^3) See the following post How do you find the derivative of this using the differentiation formula?

OpenStudy (wesdg1978):

\[g(x)=\frac{ 1 }{ \sqrt[4]{x^3} } \]

OpenStudy (blurbendy):

i would start by rewriting the denominator as x^(3/4)

OpenStudy (anonymous):

well, this is equal to \[g(x)= x ^{\frac{ -3 }{ 4 }}\] So, g'(x) would be?

OpenStudy (wesdg1978):

-3/4 x^-7/4

OpenStudy (anonymous):

Use Power Rule

OpenStudy (jhannybean):

\[\large g(x) = x^{-3/4}\]Use the power rule.\[\large g'(x) = -\cfrac{3}{4} x^{-3/4-1}\]\[\large g'(x) = -\cfrac{3}{4} x^{-7/4} \]

OpenStudy (jhannybean):

\[\large -\cfrac{3}{4x^{7/4}} \ \text{or} \ -\cfrac{3}{4\sqrt[4]{x^7}} \]

OpenStudy (anonymous):



OpenStudy (jhannybean):

Oooook.

OpenStudy (jhannybean):

Do you understand, @wesdg1978 ?

OpenStudy (wesdg1978):

@Jhannybean Yes, I understand now, thanks for your help!

OpenStudy (jhannybean):

Awesome :)